New Publications (TTN 1,1)

Claudio Beccari. L^{A}TEX — Guida a un sistema di editoria elettronica. Milano: TUGboat 13, 2. Editore Ulrico Hoepli Milano, 1991. Pp. ix + 399. ISBN 88-203-1931-4.

• The book is the first and only one in Italian at this moment. Although it contains an introductory chapter for the beginner, its main purpose is to give the reader advanced information on L^{A}TEX (and partially on TEX) in order to prepare sophisticated macros, style files, and to make clever use of fonts.

Arvind Borde. TEX by Example. New York: Academic Press, 1991. Pp. 169 TUGboat 13, 4. (softcover), US$19.95. ISBN 0-12-117650-9.

• This book teaches entirely by example. Input and output are shown sideby-side, allowing users to spot quickly how to get symbols, how to obtain italics, how to use bold type, how to start new paragraphs, and so on. The alphabetic appendix is both an index and a glossary, which systematically lists all commonly used features of TEX.

Victor Eijkhout. TEX by Topic: A TEXnician’s Reference. Wokingham, UK: TUGboat 13, 2. Addison-Wesley, 1992. Pp. viii + 307 (softcover), US$30.00. ISBN 0-201-56882-9.

• TEX by Topic is a systematic reference manual designed as a companion to tutorial guides about TEX.

Not for the novice, it is meant for users with a basic understanding of TEX who have specific questions, or who are keen to explore the full potential of this unique computer typesetting system.

Jane Hahn. L^{A}TEX for Everyone. Mill Valley: Personal TEX Inc., 1991. Pp. xi TUGboat 13, 2. + 346 (softcover), US$19.95. ISBN 0-9631044-0-3.

• L^{A}TEX for Everyone is a tutorial as well as a reference guide. Written with the beginning user in mind, it is designed to be easy to read and understand, and contains examples of all the basic commands. The experienced L^{A}TEX user, however, will also find this book to be a valuable reference tool.

Raymond Seroul and Silvio Levy. A Beginner’s Book of TEX. New York: TUGboat 13, 1. Springer Verlag, 1991. Pp. 283 (softcover), US$29.95, DM58,– . ISBN 0-387-97562-4.

• Seroul and Levy’s book is an excellent beginner’s text. It is meant for careful reading, in which case it gives a surprising amount of detail. The book also has a reference section. This is an expanded translation of Seroul’s book, Le petit Livre de TEX (Paris: InterEditions, 1989).

André Heck, ed. Desktop Publishing in Astronomy and Space Sciences. Sing- TUGboat 13, 4. apore: World Scientific Publishing Co. Pte. Ltd., 1992. Pp. xi + 240 ISBN 981-02-0915-0.

• The volume gathers together the communcations presented at the colloquium on Desktop Publishing in Astronomy and Space Sciences, held in Strasbourg, France, 1–3 October 1991, at the Louis Pasteur University.

Wynter Snow. TEX for the Beginner. Reading, Mass.: Addison-Wesley, 1992. TUGboat 13, 4. Pp. xii + 377 + 23 (index) (softcover), US$29.25. ISBN 0-201-54799-6.

• This book is a carefully paced, tutorial introduction for people first learning the [TEX] system. Special emphasis is given to what can go wrong, and how to fix it when it does. L^{A}TEX notes are provided for use with a set of macros.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: no. 12, décembre 1991, 85pp. ISSN 1140-9304. Official publication of GUTenberg, Groupe francophone des Utilisateurs de TEX.

MAPS: issue 92.1, May 1992, 162pp. Official publication of NTG, Nederlandstalige TEX Gebruikersgroep (MAPS = “Minutes and APendiceS”). Two issues per year. Contact: ntg@hearn.bitnet.

TEXbulletin: no. 1, 1992, 48pp. Official newsletter (in Czech and Slovak) of CsTUG. Previous 4 issues in 1991 appeared without numbering. Four issues per year. Contacts: Karel Horak horakk@csearn.bitnet, Jiří Veselý

jvesely@cspguk11.bitnet, or Jiří Zlatuska zlatuska@cspuni12.bitnet.

TEXline: issue 14, February 1992, 28pp. Contact Malcolm Clark for details

malcolmc@sun.pcl.ac.uk.

Die TEXnische Komödie: vol. 4, no. 1, mai 1992, 68pp. Official newsletter of DANTE, Deutschsprachige Anwendervereinigung TEX e.V. Four issues per

year. Contact: dante@dhdurz1.bitnet or dante@vm.urz.uni-
heidelberg.de.

Note: For more information on the various user groups, consult pp. 119–125 in TUG’s 1991 Resource Directory.

Jiří Zlatuška, ed. EuroTEX 92, Proceedings of the 7th European TEX Conference. Brno: Masaryk University and Prague: Czechoslovak TEX Users Group (CsTUG), 1992. ISBN 80-210-0480-0.

• Copies can be ordered directly from CsTUG for the price of DM 30.– plus postage (or the equivalent in some other currency) at the following address: Československé sdružení uZivatelů TEXu (CsTUG), Matematický ústav UK, Sokolovská 83, CS-186 00 Praha 8, Czechoslovakia. They may also be available through your local TEX users group.^{1}

Stephan von Bechtolsheim. TEX in Practice. New York: Springer Verlag, Oct. 1992. 4-volume set (hardcover), US$169.00. ISBN 0-387-97296-X. Volumes also available individually, US$49.00 each:

Vol. 1: Basics. Pp. 359, 9 illus. ISBN 0-387-97595-0.

Vol. 2: Paragraphs, Math, and Fonts. Pp. 384. ISBN 0-387-97596-9. Vol. 3: Tokens, Macros. Pp. 544, 22 illus. ISBN 0-387-97597-7. Vol. 4: Output Routines, Tables. Pp. 300. ISBN 0-387-97598-5.

• “A recent surge of good TEX implementations for PCs has put TEX on the disks of many people including writers, designers, desktop publishers, and engineers. With such increased interest in TEX, there is a need for good TEX books. TEX in Practice is the ideal reference and guide for the TEX community. The four-volume set is written by an acknowledged expert in the field and addresses the needs of the TEX novice to the more experienced ‘TEXpert.’ The book provides step-by-step introduction to the various functions of TEX with many relevant examples.”

Stefan Schwartz and Rudolf Potuček. L^{A}TEX — Satzkunst statt DTP. [“L^{A}TEX — The Art of Typography rather than DTP”]. (Special issue in the series CHIP-Anwender-Praxis.) 1992. DM 34,– . ISBN 3-8023-1178-7.

• Can be obtained from: Vogel Verlag, Postfach 67 40, D-8700 Würzburg, Germany.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: no. 13 juin 1992, 90pp. ISSN 1140-9304.

1The Table of Contents of the Proceedings appears on p. 28 of this issue, and a short report on the conference by Alan Hoenig appears on pp. 22–23.

Rames Abdelhamid. Das Vieweg L^{A}TEX-Buch: Eine praxisorientierte Einführung. Braunschweig: Vieweg-Verlag, 1992. About 200pp. ISBN 3-528-05145-0.

• This book provides a well-structured and readable introduction to L^{A}TEX in combination with the german.sty option.

George Gratzer. Math into TEX. New York: Springer Verlag. Pp. 187 (softcover). Approx. $34.00 US. ISBN 0-8176-3637-4. [Avail. Dec. 1992]

• “This book is for the mathematician, engineer, or scientist who wants to write and typeset articles with mathematical formulas but who does not want to spend a great deal of time learning how to do it. It assumes little familiarity with TEX or L^{A}TEX.” [From promotional material]

Piotr Wyrostek, trans. L^{A}TEX: System przygotowywania dokumentöw. Przewodnik ürytkownika i prodrëcznik. Polish translation of Lamport’s L^{A}TEX manual. Krakow: Ariel. Price: 91000 zl. ISBN 83-900460-1-6.

Michaela Lichá and Oldřich Ulrych. AMS-TEX verze 2.1. Prague: Czechoslovak TEX Users Group, 1992. Pp. 80 [In Czech]

• “This publication provides a short (but complete) tour through AMS-TEX 2.1 together with amsppt.sty 2.1. Almost all features are demonstrated with a lot of examples. The book is organized in such a way that it is useful for novices (basic textbook), and for advanced users (quick reference quide) as well. Contains a table of all symbols.” [Info. from authors]

Latest issues of other TEX newsletters

Baskerville: vol. 2, no. 1, March 1992, 12pp. Official newsletter of the UK TEX

Users’ Group. Die TEXnische Komödie: vol. 4, no. 2, August 1992, 64pp. Official newsletter of

DANTE, Deutschsprachige Anwendervereinigung TEX e.V. MAPS: issue 92.2, 1992, 176pp. Official publication of NTG, Nederlandstalige

TEX Gebruikersgroep (MAPS = “Minutes and APendiceS”). TEXbulletin: no. 2, 1992, 48pp; no. 3, 1992, 64pp. Official newsletter of CsTUG.^{2} TUGboat: 13, no. 3, 1992, 162pp. Proceedings of TUG’92 meeting in Portland,

Oregon.

2Re: TTN 1,3 notice on the Prague Proceedings: Jiří Zlatuska of CsTUG asks that orders for the publication include a shipping and handling charge: $5 for surface mail, $10 for airmail in addition to the 30,– DM cost. –Ed.

Donald E. Knuth. Literate Programming. (CSLI Lecture Notes 27.) Stanford: Center for the Study of Language and Information, 1992. Pp. xiii + 368 £19.95 (UK). ISBN 0-937073-80-6 (paper), ISBN 0-937073-81-4 (cloth).

• ‘Literate Programming’ (using the WEB system) was developed by Knuth and used to write TEX and its documentation. This book is a collection of previously published articles and extracts from books, together with additional references and some other new material. In particular, it contains the papers “The Errors of TEX” and the “Error Log for TEX, 1978–91”.

Michael Vulis. Modern TEX and Its Applications. Boca Raton: CRC Publish- TUGboat 14, 1. ers, 1992. Pp. 275, includes diskette. US$32.95 ($39.95 outside the US). ISBN 0-8493-4431-X (paper).

• “A guide to plain TEX (and VTEX), intended for scientists, etc., and support staff who prepare technical texts and documents. It covers all basic topics for understanding the TEX user environment, including fonts and characters, formatting, math mode, macros, terminal and file operations, tables, foreign language capabilities (including Cyrillic), indices, references, tables of contents, scalable typefaces, and graphics. A PC-compatible disk containing examples, extra typefaces, and a ready-to-run restricted [!] version of TEX is included.” [adapted from promotional material]

Antoni Diller. L^{A}TEX Line by Line: Tips and Techniques for Document Pro- TUGboat 14, 2. cessing. (Software Engineering.) Chichester: John Wiley and Sons, 1992. Pp. 200 (softcover). US$29.95. ISBN 0-471-93471-2.

• “A handbook to the L^{A}TEX text processing system, this easy-to-follow guide teaches all users how to produce a wide variety of documents — from business letters to technical reports. The book offers ready-to-use templates for producing articles, reports, letters and more; explains how to quickly generate tables of contents, indices and bibliographies; and includes many practical examples, an appendix of mathematical symbols and a complete glossary of typesetting terms.” [adapted from promotional material]

G. Grätzer. Math into TEX: A Simplified Introduction Using AMS-L^{A}TEX.

Basel: Birkhäuser, 1992. SFr 52.–, DM 57.–. ISBN 3-7643-3637-4. Note: This item was already listed in the last issue. This entry adds the bibliographic data of the co-publisher.

Latest issues of other TEX newsletters

Die TEXnische Komödie: vol. 4, no. 3, November 1992, 56pp (DANTE).

Malcolm Clark. A plain TEX Primer. Oxford: Oxford University Press, 1992. Pp. 481. ISBN 0-19-853784-0 (Hbk), ISBN 0-19-853724-7 (Pbk).

• “This primer introduces plain TEX, and aims to help the reader get started. It explains why TEX approaches its subject in the way it does, and provides the context into which it fits. Another thread is the emphasis placed on document structure. Stress is also laid on practical work. The book is aimed at scientists and researchers preparing their own books and papers, or technical typists used to the conventions and jargon of their field.” [adapted from the publisher’s description]

Stanley A. Sawyer & Steven Krantz. A TEX Primer for Scientists. Boca Raton: CRC Publishers, 1993. Pp. 175. US$29.95 ($36.00 outside USA). ISBN 0-8493-7159-7.

• “This concise, straightforward text covers all basic topics of equation typesetting, and some advanced topics: how to automatically generate a table of contents, and indices; how to make text flow around a graphic; how to add more fonts; how to change fonts in math mode (for titles, etc.). In an appendix differences between TEX, L^{A}TEX and AMS-TEX are discussed.” [adapted from promotional material]

Hans Ibrus. Väike Eesti TEXiraamat. [ The Small Estonia TEXbook ]. Tartumaa, Estonia: a/S “KompuT” kirjastus, 1992. Pp. 144.

• Preface by Enn Saar. [Further description on p. 18.] Appendix A is a translation into Estonian of First Grade TEX, by Arthur Samuel.

Gordon Müllejans. METAFONT. Eine Referenz. Bonn: Addison-Wesley, 1992. Pp. 176. DM 39,90. ISBN 3-89319-363-4.

• This is neither a tutorial nor a textbook on . It is a reference tool, and consists mainly of listings with short descriptions, in particular of the basic commands, the macros of cmbase, and the parameters used to define the CM fonts.

Latest issues of other TEX newsletters

Die TEXnische Komödie: vol. 4, no. 4, Dezember 1992, 56pp.

La Lettre GUTenberg: numéro 0, février 1993, 16pp.

Helmut Kopka and Patrick W. Daly. A Guide to L^{A}TEX: Document Prepa- TUGboat 14, 2. ration for Beginners and Advanced Users. Addison-Wesley (UK) Ltd., 1993. Pp. 448. US$34.50. ISBN 0-201-56889-6.

• This book is based on two well-known German books by Helmut Kopka: L^{A}TEX — Eine Einführung, and L^{A}TEX — Erweiterungsmöglichkeiten.

Francesc Rosselló Llompart. L’ABC del TEX. Palma de Mallorca: UIB-DMI. Departament de Ciències Matemàtiques i Informàtica, Universitat de les Illes Balears, 1991.

Reinhard Wonneberger. L^{A}TEX. 3., durchgesehene und erweiterte Auflage.

(Addison-Wesley Kompaktführer). Bonn, etc.: Addison-Wesley, 1993. Pp. xvi + 166. DM 29,80. ISBN 3-89319-589-0.

• This is the third (expanded) edition of a concise German L^{A}TEX reference (previous editions 1987 and 1988). The main addition is a section on german.sty, which is widely used in German-speaking countries.

Electronic Publications

Hubert Partl, Elis. Schlegl and Irene Hyna. Una Descripción de L^{A}TEX.

• Spanish translation by Tomas Bautista of well-known German introduction to L^{A}TEX, L^{A}TEX-Kurzbeschreibung. Includes material on the spanish style option. Available in Spain from ftp.eunet.es (also known as goya.

uu.es) in the directory info/unix/text/TeX\spanish/tutorial/, as well

as from the main TEX archives.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: No. 14: EuroTEX’92: Proceedings of the 7th European TEX Conference. Prague, Sept. 14–18, 1992. Pp. vi + 330. No. 15, avril 1993, 61pp. ISSN 1140-9304.

GUST: Zyszyt 1 (1993). [March issue 35 + 7; A4 format.] • First issue of the newsletter GUST from the Polish user group Grupa Użytkowników Systemu TEX. Editors are: Włodek Bzyl and Tomasz Przechlewski. Address is: Instytut Matematyki UG, Wita Stwosza 56, 80-952 Gdańsk. E-mail:

matwb@halina.univ.gda.pl.

MAPs: issue 93.1, 1993, 233pp. [Dutch-speaking user group, NTG publication].

Arvind Borde and Tomas Rokicki: A User’s Guide for TEXHelp; The On-Line TEX Handbook. Academic Press, 1993. v + 28pp. ISBN 0-12-1176040-1. • This small booklet documents the TEXHelp program for IBM PC and compatible microcomputers.

Arvind Borde: Mathematical TEX by Example. Academic Press, 1993. 356pp. US$19.95. ISBN 0-12-117645-2.

• As the title suggests, this book is an introduction to TEX by means of examples. They almost entirely consist of pairs of pages: one contains some piece of TEX code (with explanations and comments added via footnotes), while the other shows the output produced by this code.

Nicholas J. Higham: Handbook of Writing for the Mathematical Sciences. SIAM — Society for Industrial and Applied Mathematics, Philadelphia, 1992. xii + 241pp. US$21.50. ISBN 0-89871-314-5.

• This book is intended for all authors of technical papers (and books) or talks, and covers virtually all relevant issues. The author discusses questions of style as well as practical matters. The use of computers in the publishing process — including TEX, of course — is discussed in detail. An appendix contains a list of TEX and L^{A}TEX symbols. Among the various useful materials are hints for Internet use and an extensive bibliography.

Christian Rolland: L^{A}TEX guide pratique. Addison-Wesley France, July 1993. 280pp. approx. 190 FF. ISBN 2-87908-025-0. (Distributed by Bordas, Bordas code 009808.)

Newsletters

GUST: Zeszyt 2 (1993), 42pp. Official newsletter of the Polish User Group, Grupa Użytkowników Systemu TEX.

La Lettre GUTenberg: no. 1, juillet 1993, 16pp.

Die TEXnische Komödie: 5. Jahrgang, Heft 1, Mai 1993, 72pp; Heft 2, Sep. 1993.

TUGboat: vol. 14, no. 1, April 1993, 92pp; vol. 14, no. 2, July 1993, 64pp.

Other books

Paul Abrahams. TEX for the Impatient. Reading, Mass.: Addison-Wesley, TUGboat 11, 4. 1990. 357 pp (softcover).

• “TEX for the Impatient addresses the practical demands of busy people who have discovered the unmatched benefits of TEX, but whose time and interest in mastering all its intricacies may be limited … anyolne … who needs a quick reference guide to TEX.”

Wolfgang Appelt. TEX für Fortgeschrittene. [TEX for the Advanced]. Bonn: TUGboat 11, 2. Addison-Wesley, 1988. ISBN 3-89319-115-1.

Arvind Borde. Mathematical TEX by Example. Academic Press, 1993. ISBN TUGboat 14, 1. 0-12-117645-2. 356pp.

David J. Buerger. L^{A}TEX for Engineers and Scientists. McGraw-Hill, 1990. TUGboat 12, 2. ISBN 0-07-008845-4. 198pp (softcover).

Helmut Kopka. L^{A}TEX — Eine Einführung. 3rd rev. ed. Bonn: Addison- TUGboat 13, 1. Wesley, 1991. 375pp (hardcover).

Helmut Kopka. L^{A}TEX — Erweiterungsmöglichkeiten mit einter Einführung in . 2nd rev. ed. Bonn: Addison-Wesley, 1991. 552pp. (hard- TUGboat 13, 1. cover).

Norbert Schwartz. Einführung in TEX. Bonn, Addison-Wesley, 1988. ISBN 3-925118-97-7.

Norbert Schwartz. Inleiding TEX. Addison-Wesley Europe, Amsterdam, 1990.

ISBN 90-6789-151-7.

Norbert Schwartz. Introduction to TEX. Reading, Mass.: Addison-Wesley, TUGboat 11, 2. 1990. 278 pp (paperback). [Based on 2nd ed. of German original Einführung in TEX.]

Michael Urban. Premiers pas en L^{A}TEX. Adaptation française de Éric Cornelis. GUTenberg, 1991. 87pp (softcover).